### v1

# INFusion: Diffusion Regularized Implicit Neural Representations for 2D and 3D accelerated MRI reconstruction

Abstract
Implicit Neural Representations (INRs) are a learning-based approach to accelerate Magnetic Resonance Imaging (MRI) acquisitions, particularly in scan-specific settings when only data from the under-sampled scan itself are available. Previous work demonstrates that INRs improve rapid MRI through inherent regularization imposed by neural network architectures. Typically parameterized by fully-connected neural networks, INRs support continuous image representations by taking a physical coordinate location as input and outputting the intensity at that coordinate. Previous work has applied unlearned regularization priors during INR training and have been limited to 2D or low-resolution 3D acquisitions. Meanwhile, diffusion-based generative models have received recent attention as they learn powerful image priors decoupled from the measurement model. This work proposes INFusion, a technique that regularizes the optimization of INRs from under-sampled MR measurements with pre-trained diffusion models for improved image reconstruction. In addition, we propose a hybrid 3D approach with our diffusion regularization that enables INR application on large-scale 3D MR datasets. 2D experiments demonstrate improved INR training with our proposed diffusion regularization, and 3D experiments demonstrate feasibility of INR training with diffusion regularization on 3D matrix sizes of $256\times 256\times 80$.
I Introduction
Magnetic Resonance Imaging (MRI) suffers from lengthy acquisition times, but sampling the acquired Fourier (k-space) data below the Nyquist rate and solving the ill-posed reconstruction problem accelerates acquisitions, reducing costs to patients. Widespread reconstruction algorithms include parallel imaging [1–3] and compressed sensing [4], but deep learning based techniques exploiting training databases yield state of the art results [5–10], with diffusion models receiving recent attention as they decouple the measurement model and learned prior.
The machinery of Implicit Neural Representations (INRs) [11,12] serves as an alternative learning-based approach for accelerating MRI, particularly in scan-specific settings where only data from the under-sampled scan itself are available. Previous INR work leverages traditional regularization [13], the inherent regularization of the dense neural network architecture and positional encoding [12,14–17], and periodic activations [11]. However, previous work applies unlearned regularization priors when applying INRs for MRI reconstruction and are mainly limited to 2D or low-resolution 3D acquisitions [13–17].
Drawing inspiration from recent work in computer vision that uses diffusion priors to reconstruct INR-representations of 3D scenes from few camera views [18], we present INFusion, a technique to regularize the optimization of INRs from under-sampled MR measurements with pre-trained diffusion models for improved image reconstruction. In addition, we propose a hybrid 3D approach with our diffusion regularization that enables INR application on large-scale 3D MR datasets.
II Methods
II-A MRI Inverse Problems with INRs
The model-based accelerated MRI inverse problem reads:
$\arg\min_{x}||y-Ax||_{2}^{2}+\lambda R(x),$ (1)
where $y\in\mathbb{C}^{M}$ are the discrete k-space measurements, $A\in\mathbb{C}^{M\times N}$ is the linear measurement model, $x\in\mathbb{C}^{N}$ is the discretized image to reconstruct, and $\lambda R(\cdot):\mathbb{C}^{N}\rightarrow\mathbb{R}$ imposes regularization (e.g., sparsity, low-rank).
Let $I_{\theta}:\mathbb{R}^{C}\rightarrow\mathbb{C}$ be an INR parameterized by weights $\theta$, which represents the image as a function of the continuous coordinate domain. Then, the following optimization problem estimates weights of the INR through comparison to the acquired discrete k-space measurements [13–17]:
$\theta^{*}=\arg\min_{\theta}||y-AI_{\theta}(r)||_{2}^{2}+\lambda R(I_{\theta}(r)),$ (2)
where $r\in\mathbb{R}^{N\times C}$ indexes the spatial coordinates of the discretized voxel locations of $x$. Evaluating $I_{\theta^{*}}(r)$ then yields the desired image.
II-B INFusion: Regularizing INRs with Diffusion Models
Our proposed INFusion method regularizes the INR MRI inverse problem with pre-trained generative diffusion models as they are powerful priors that decouple from the measurement model [8-10] enabling applications with INRs. As it is computationally expensive to evaluate the diffusion model at every 3D coordinate, we also propose a stochastic regularizer in which the diffusion model is queried at random spatial coordinates $z\subset r$ in each iteration.
During each iteration, $j$, of optimization, we propose the following procedure:
1:Compute data consistency loss, $L_{data}=||y-AI_{\theta_{j}}(r)||_{2}^{2}$.
2:Select spatial coordinates $z$ so that evaluating $I_{\theta_{j}}(z)$ yields random discretized 2D slices from the current estimate of image volume $x$.
3:Compute $q_{j}=I_{\theta_{j}}(z)+n$ where $n\sim\mathcal{N}(0,\sigma^{2})$ and $\sigma\sim\mathcal{U}(\sigma_{min},1)$.
4:Run diffusion [19,20] with initialization $q_{j}$ to give a prior sample $d_{j}$.
5:Compute $L_{diffusion}=LPIPS(I_{\theta_{j}}(z),d_{j})$ with LPIPS perceptual loss [21].
6:Compute $L_{total}=L_{data}+w(j)L_{diffusion}$ with adaptive scalar weighting $w(j)$, and backpropagate to update $\theta_{j}$.
Fig. 1 presents a schematic summary of the proposed procedure.
Figure 1: Our proposed approach, INFusion, first computes standard data consistency loss $L_{data}$ by evaluating the INR at discrete voxel locations, applying the MRI forward model to the resultant image, and then comparing to the acquired k-space data. A second loss $L_{diffusion}$ is computed by adding gaussian noise to random slices produced by the INR, running prior sampling with the diffusion model, and comparing the resultant prior sample to the INR produced random slices with perceptual loss. The diffusion loss guides the INR to produce images that better match the learned prior.
III Experiments
III-A In-vivo 2D Single- and Multi- Coil Brain
We used axial, T2-weighted, single-coil and 4-coil 2D k-space from the fastMRI [22] dataset, cropped both to $192\times 192$ resolution, and retrospectively under-sampled with a 2D-Poisson pattern generated with BART [23] corresponding to an acceleration rate of $R=4$ for single-coil ($R$ times less than Nyquist) and $R=6$ for multi-coil. Both experiments compared the discretized image produced by solving the standard MRI inverse problem with L1-Wavelet regularization [24] to an INR trained with (i) none, (ii) L1-Wavelet, or (iii) our proposed diffusion regularization.
Second, we took 96 additional, multi-coil 2D samples from the fastMRI dataset, and retrospectively under-sampled them with 2D-Poisson under-sampling masks with an acceleration rate of $R=8$ and $R=9$. These samples had coil counts ranging from $4-20$. We compared error of the discretized images produced by solving the standard MRI inverse problem with L1-Wavelet regularization and INRs with L1-Wavelet and our proposed diffusion regularization.
III-B In-vivo small 3D Single Coil Brain and large 3D Multi-coil Knee
We used a multi-slice single-coil $192\times 192\times 16$ dataset from fastMRI with axial brain slices, and treated it as 3D by retrospectively under-sampling it in the $y-z$ direction with a $R=2$ Poisson under-sampling mask.
Finally, we used k-space of an 8-coil 3D knee volume from the SKM-TEA [25] dataset, resized it to $256\times 256\times 80$, and uniformly under-sampled by $R=4\times 1$ in the $y-z$ directions. Fitting the problem in GPU memory and solving it in feasible time required two changes to our INR procedure:
  • Only the $x-y$ dimensions were encoded with coordinates, and the model instead returned 80 discretized slice values at each coordinate.
  • Diffusion regularization was computed with respect to two random slices each iteration.
Both 3D experiments compared discretized images produced by an INR trained with and without our INFusion method. Note that our diffusion model was trained on $x-y$ images, but under-sampling was performed in the $y-z$ direction.
III-C INR Details
All INRs consisted of fully-connected neural-networks with ReLU activations and 128 gaussian Fourier-features [12]. 2D, 3D brain, and 3D knee networks had 165K, 1643K, and 21411K parameters, respectively.
III-D Diffusion Model Details
We trained separate diffusion models, using an architecture and training procedure provided in [20], for the brain model with 8500 axial fastMRI slices and the knee with 15720 sagittal SKM-TEA slices.
IV Results
Figure 2: Quantitative comparison of standard L1-Wavelet CS and INRs trained with Wavelet and our proposed regularization at $R=8$ and $R=9$ across 96 samples. Our proposed INFusion approach improves NRMSE across the dataset.
INFusion achieved lower normalized-root-mean-squared-error (NRMSE) on the single (Fig. 3, A) and multi (Fig. 3, B) coil brain slice in comparison to INRs trained with Wavelet or without regularization and the standard Compressed Sensing (CS) L1-Wavelet MRI inverse problem. The box plot in Fig. 2 shows that INFusion improves reconstruction performance with respect NRMSE over 96 slices in comparison to standard CS L1-Wavelet and INRs with Wavelet regularization.
Figure 3: Reconstructions of discrete images on the (A.) 2D single-coil brain data at $R=4$ and (B.) 2D multi-coil brain data at $R=6$ with standard L1-wavelet Compressed Sensing (CS) and INRs with none, L1-wavelet, and our proposed diffusion regularization. The proposed INFusion approach yields images with lowest NRMSE.
In Fig. 4, A, INFusion outperforms no regularization on the modest, single-coil 3D dataset, but encoding all three spatial dimensions in the INR required prohibitively large GPU memory usage and compute. Fig. 4, B shows that our proposed approach of diffusion regularization with random slices and just encoding the $x-y$ dimensions in coordinates enables training of INRs on $256\times 256\times 80$ matrix size k-space.
Figure 4: (A.) INFusion outperforms no regularization on the modest, single-coil 3D dataset, but encoding all three spatial dimensions in the INR required prohibitively large GPU memory usage and compute. (B.) Applying diffusion regularization to random slices at each iteration and encoding only the x-y coordinates enables training of INRs with diffusion regularization on a realistically sized $256\times 256\times 80$ 3D k-space volume.
V Discussion
INFusion exploits the generative capabilities of diffusion models to regularize INR training from scan-specific MR measurements. We show improved performance in 2D and enable application in realistic 3D settings where the diffusion models were also trained on a different orientation than the under-sampling. In this way, we regularize 3D reconstructions without an explicit 3D prior.
Future work will investigate the effectiveness of INRs as an image representation and explore the trade-offs between using diffusion models as regularization for INRs versus using them for posterior sampling directly on the discretized voxel grid.
Acknowledgment
This work was supported in part by NSF CCF-2239687 (CAREER), NSF IFML 2019844, and JCCO postdoctoral fellowship.

### v2

@@
 I Introduction
-Magnetic Resonance Imaging (MRI) suffers from lengthy acquisition times, but sampling the acquired Fourier (k-space) data below the Nyquist rate and solving the ill-posed reconstruction problem accelerates acquisitions, reducing costs to patients.
-Widespread reconstruction algorithms include parallel imaging [1–3] and compressed sensing [4], but deep learning based techniques exploiting training databases yield state of the art results [5–10], with diffusion models receiving recent attention as they decouple the measurement model and learned prior.
-The machinery of Implicit Neural Representations (INRs) [11,12] serves as an alternative learning-based approach for accelerating MRI, particularly in scan-specific settings where only data from the under-sampled scan itself are available.
-Previous INR work leverages traditional regularization [13], the inherent regularization of the dense neural network architecture and positional encoding [12,14–17], and periodic activations [11].
-However, previous work applies unlearned regularization priors when applying INRs for MRI reconstruction and are mainly limited to 2D or low-resolution 3D acquisitions [13–17].
-Drawing inspiration from recent work in computer vision that uses diffusion priors to reconstruct INR-representations of 3D scenes from few camera views [18], we present INFusion, a technique to regularize the optimization of INRs from under-sampled MR measurements with pre-trained diffusion models for improved image reconstruction.
-In addition, we propose a hybrid 3D approach with our diffusion regularization that enables INR application on large-scale 3D MR datasets.
+Magnetic Resonance Imaging (MRI) is a clinically important medical imaging modality that provides excellent soft tissue contrast without any ionizing radiation.
+However, MRI acquisitions suffer from an inherent trade-off between scan time, resolution, and signal-to-noise ratio (SNR) [1].
+Shortening MRI scan time while maintaining image quality through computation reduces costs to patients and hospitals and reduces sensitivity to motion artifacts.
+MRI scans can be accelerated by sampling the acquired Fourier (k-space) data below the Nyquist sampling rate and solving the ill-posed reconstruction problem to produce a high fidelity image from under-sampled measurements.
+To reduce scan time, clinics routinely use parallel imaging, [2, 3, 4], which interpolates missing data with the encoding capability of the multi-coil signal receive array, and compressed sensing [5], which exploits random sampling and sparsity.
+Recently, machine learning algorithms for accelerated MRI reconstruction employing training databases yield state of the art results [6, 7, 8, 9, 10, 11, 12].
+End-to-end methods learn a point-wise mapping between under-sampled and fully-sampled data but are susceptible to test time shifts in the measurement operator.
+More recently, generative methods have been proposed that learn a prior over clean images and are robust to test time shifts in the forward operator.
+All the MRI reconstruction algorithms previously discussed assume that a discretized grid of voxels represents the desired MRI image.
+However, alternative representations can also be employed, such as Implicit Neural Representations (INRs) [13, 14].
+An INR trains a neural network to take as input a spatial coordinate and output the signal value of interest at that spatial coordinate, thus providing a continuous image representation, and in some sense, implicitly storing the image in the weights of the neural network.
+Recent work utilizes INRs for accelerating MRI, particularly in scan-specific settings where only data from the under-sampled scan itself are available.
+By combining traditional regularization with the inherent regularization of the dense neural network architecture, positional encoding, and periodic activations, previous work demonstrates that INRs help further accelerate MRI scans [15, 16, 17, 18, 19].
+However, these methods apply unlearned spatial regularization priors and are mainly limited to 2D or low-resolution 3D acquisitions.
+This work proposes further accelerating MRI with INR-based reconstructions by drawing inspiration from recent work in computer vision that uses diffusion priors to reconstruct INR-representations of 3D scenes from few camera views [20].
+In particular, our method, termed INFusion, uses diffusion models pre-trained on MRI images to regularize the optimization of INRs from under-sampled measurements.
+First, we demonstrate that our proposed approach improves reconstruction performance in comparison to competing techniques in 2D, $T_{2}$-weighted brain MRI.
+Then, we propose a hybrid 3D approach with our diffusion regularization that enables INR application on large-scale 3D MR datasets with 2D diffusion models, and we show experiments on a realistic, 3D knee acquisition.
 II Methods
-II-A MRI Inverse Problems with INRs
-The model-based accelerated MRI inverse problem reads:
+II-A The Accelerated MRI Inverse Problem
+Given the linear, multi-channel MRI measurement model $A=PFS\in\mathbb{C}^{(MC)\times N}$, solving the following inverse problem yields an image from under-sampled measurements, $y\in\mathbb{C}^{MC}$, acquired in an accelerated scan,
 $\arg\min_{x}||y-Ax||_{2}^{2}+\lambda R(x),$ (1)
-where $y\in\mathbb{C}^{M}$ are the discrete k-space measurements, $A\in\mathbb{C}^{M\times N}$ is the linear measurement model, $x\in\mathbb{C}^{N}$ is the discretized image to reconstruct, and $\lambda R(\cdot):\mathbb{C}^{N}\rightarrow\mathbb{R}$ imposes regularization (e.g., sparsity, low-rank).
-Let $I_{\theta}:\mathbb{R}^{C}\rightarrow\mathbb{C}$ be an INR parameterized by weights $\theta$, which represents the image as a function of the continuous coordinate domain.
-Then, the following optimization problem estimates weights of the INR through comparison to the acquired discrete k-space measurements [13–17]:
-$\theta^{*}=\arg\min_{\theta}||y-AI_{\theta}(r)||_{2}^{2}+\lambda R(I_{\theta}(r)),$ (2)
-where $r\in\mathbb{R}^{N\times C}$ indexes the spatial coordinates of the discretized voxel locations of $x$. Evaluating $I_{\theta^{*}}(r)$ then yields the desired image.
-II-B INFusion: Regularizing INRs with Diffusion Models
-Our proposed INFusion method regularizes the INR MRI inverse problem with pre-trained generative diffusion models as they are powerful priors that decouple from the measurement model [8-10] enabling applications with INRs.
+where $x\in\mathbb{C}^{N}$ is the discretized image to reconstruct, $S\in\mathbb{C}^{(NC)\times N}$ describes how the multi-coil receive array measures signal, $F\in\mathbb{C}^{(NC)\times(NC)}$ applies the Fourier Transform to each coil image, $P\in\mathbb{C}^{(MC)\times(NC)}$ indicates which k-space points are sampled (with $M<N$ for reduced scan time), and $\lambda R(\cdot):\mathbb{C}^{N}\rightarrow\mathbb{R}$ imposes regularization (e.g., sparsity, low-rank).
+II-B Implicit Neural Representations
+Let $r\in\mathbb{R}^{3}$ represent a physical coordinate corresponding to a spatial location of the MRI anatomy to image (without loss of generality assume that $r\in[0,1]^{3}$).
+Then, an Implicit Neural Representation (INR) is a fully connected neural network $I_{\theta}:\mathbb{R}^{3}\rightarrow\mathbb{C}$ that takes a spatial coordinate, $r$, as input and outputs the signal value of interest of the MRI anatomy at that spatial location.
+Given a discretized image $x$, the INR can be trained by optimizing
+$\theta^{*}=\arg\min_{\theta}||x-I_{\theta}(\boldsymbol{r})||_{2}^{2},$ (2)
+where $\boldsymbol{r}\in\mathbb{R}^{3\times N}$ is the spatial coordinates of the voxel locations in the discretized image.
+Then, $I_{\theta^{*}}$ can be evaluated at spatial coordinates not on the original grid of coordinates, $\boldsymbol{r}$, thus yielding a continuous representation.
+II-C Accelerated MRI with INRs
+Given the discrete, under-sampled k-space measurements $y$, the following optimization problem estimates weights of an INR in the accelerated MRI setting:
+$\theta^{*}=\arg\min_{\theta}||y-AI_{\theta}(\boldsymbol{r})||_{2}^{2}+\lambda R(I_{\theta}(\boldsymbol{r})),$ (3)
+recalling that $\boldsymbol{r}$ is the spatial coordinates of the discretized voxel locations.
+Then evaluating $I_{\theta^{*}}(\boldsymbol{r})$ yields the desired image.
+Previous work shows that the structure of the INR helps regularize the ill-posed reconstruction problem and yields higher fidelity images in comparison to directly solving for the discretized image [15, 16, 17, 18, 19].
+However, previous work regularizes the optimization problem with a hand-crafted spatial prior $R$, leaving room for further improvements with learned regularizors.
+II-D INFusion: Regularizing INRs with Diffusion Models
+Our proposed INFusion method regularizes the INR MRI inverse problem with pre-trained generative diffusion models as they are powerful priors that decouple from the measurement model [9, 10, 11] enabling applications with INRs.
 As it is computationally expensive to evaluate the diffusion model at every 3D coordinate, we also propose a stochastic regularizer in which the diffusion model is queried at random spatial coordinates $z\subset r$ in each iteration.
@@
 2:Select spatial coordinates $z$ so that evaluating $I_{\theta_{j}}(z)$ yields random discretized 2D slices from the current estimate of image volume $x$.
-3:Compute $q_{j}=I_{\theta_{j}}(z)+n$ where $n\sim\mathcal{N}(0,\sigma^{2})$ and $\sigma\sim\mathcal{U}(\sigma_{min},1)$.
+3:Compute $q_{j}=I_{\theta_{j}}(z)+n$ where $n\sim\mathcal{N}(0,\sigma^{2})$ and $\sigma\sim\mathcal{U}(\sigma_{\min},1)$.
 4:Run diffusion [21, 22] with initialization $q_{j}$ to give a prior sample $d_{j}$.
@@
 6:Compute $L_{total}=L_{data}+w(j)L_{diffusion}$ with adaptive scalar weighting $w(j)$, and backpropagate to update $\theta_{j}$.
-Fig. 1 presents a schematic summary of the proposed procedure.
+Fig.
+1 presents a schematic summary of the proposed procedure.
+In 3D settings, evaluating the INR at all spatial coordinates and backpropagating through the prior sampling procedure becomes computationally prohibitive.
+For 3D problems, we propose only encoding the $x-y$ dimension with spatial coordinates and having the model return discretized values corresponding to the third $z$ dimension.
+In addition, we compute our proposed diffusion regularization with respect to two random slices at each iteration.
 Figure 1: Our proposed approach, INFusion, first computes standard data consistency loss $L_{data}$ by evaluating the INR at discrete voxel locations, applying the MRI forward model to the resultant image, and then comparing to the acquired k-space data.
@@
 The diffusion loss guides the INR to produce images that better match the learned prior.
+II-E INR and INFusion Hyperparameters
+For all experiments, our INRs consisted of fully-connected neural networks with relu activations and 4 hidden layers using spatial coordinates encoded with 128 gaussian fourier features [14].
+INRs for 2D and 3D experiments had size 256 and size 3200 hidden layers respectively.
+For the INFusion procedure, we selected $\sigma_{\min}$ to decrease from [1,0] linearly as function of the current optimization iteration $j$.
+Similarly, $w(j)$ decreased exponentially from [1,0].
+We ran 10 iterations of the diffusion sampling formulation in [22] to generate the prior sample $d_{j}$.
+II-F Generative Diffusion Model Training
+We trained two diffusion models using a U-Net style architecture with 65 million parameters and used the ”EDM” hyper-parameters, loss function, and diffusion noise schedule [22].
+The first model was trained on 8500 $T_{2}$ axial brain slices from fastMRI [24] of dimension $192\times 192$ and the second was trained on 15720 sagittal knee slices from SKM-TEA [25].
 III Experiments
@@
 We used a multi-slice single-coil $192\times 192\times 16$ dataset from fastMRI with axial brain slices, and treated it as 3D by retrospectively under-sampling it in the $y-z$ direction with a $R=2$ Poisson under-sampling mask.
-Finally, we used k-space of an 8-coil 3D knee volume from the SKM-TEA [25] dataset, resized it to $256\times 256\times 80$, and uniformly under-sampled by $R=4\times 1$ in the $y-z$ directions.
-Fitting the problem in GPU memory and solving it in feasible time required two changes to our INR procedure:
-  • Only the $x-y$ dimensions were encoded with coordinates, and the model instead returned 80 discretized slice values at each coordinate.
-  • Diffusion regularization was computed with respect to two random slices each iteration.
+We also used k-space of an 8-coil 3D knee volume from the SKM-TEA [25] dataset, resized it to $256\times 256\times 80$, and uniformly under-sampled by $R=4\times 1$ in the $y-z$ directions.
 Both 3D experiments compared discretized images produced by an INR trained with and without our INFusion method.
+The smaller 3D brain experiment encoded all three spatial dimensions with coordinates and computed diffusion regularization across all 16 slices each iteration.
+The larger knee experiment incorporated the previously discussed modifications for 3D computational feasibility.
 Note that our diffusion model was trained on $x-y$ images, but under-sampling was performed in the $y-z$ direction.
-III-C INR Details
-All INRs consisted of fully-connected neural-networks with ReLU activations and 128 gaussian Fourier-features [12].
-2D, 3D brain, and 3D knee networks had 165K, 1643K, and 21411K parameters, respectively.
-III-D Diffusion Model Details
-We trained separate diffusion models, using an architecture and training procedure provided in [20], for the brain model with 8500 axial fastMRI slices and the knee with 15720 sagittal SKM-TEA slices.
 IV Results
 Figure 2: Quantitative comparison of standard L1-Wavelet CS and INRs trained with Wavelet and our proposed regularization at $R=8$ and $R=9$ across 96 samples.
-Our proposed INFusion approach improves NRMSE across the dataset.
+Our proposed INFusion approach improves NRMSE across the test dataset.
 INFusion achieved lower normalized-root-mean-squared-error (NRMSE) on the single (Fig.
@@
 Figure 3: Reconstructions of discrete images on the (A.) 2D single-coil brain data at $R=4$ and (B.) 2D multi-coil brain data at $R=6$ with standard L1-wavelet Compressed Sensing (CS) and INRs with none, L1-wavelet, and our proposed diffusion regularization.
-The proposed INFusion approach yields images with lowest NRMSE.
+The proposed INFusion approach yields images with lowest NRMSE and best qualitative image quality.
 In Fig.
@@
 V Discussion
-INFusion exploits the generative capabilities of diffusion models to regularize INR training from scan-specific MR measurements.
+INFusion exploits the generative capabilities of diffusion models to regularize INR training from scan-specific, under-sampled MR measurements.
+INR-based accelerated MRI inverse problems solved with our proposed diffusion regularization yield higher fidelity images with reduced scan time.
 We show improved performance in 2D and enable application in realistic 3D settings where the diffusion models were also trained on a different orientation than the under-sampling.
-In this way, we regularize 3D reconstructions without an explicit 3D prior.
-Future work will investigate the effectiveness of INRs as an image representation and explore the trade-offs between using diffusion models as regularization for INRs versus using them for posterior sampling directly on the discretized voxel grid.
+Full 3D diffusion models are challenging to train, due to computational burden and limited data availability, but INFusion enables regularization of 3D reconstruction problems using 2D diffusion models.
+In addition, the 2D diffusion models can be trained on data from any 2D orientation that is available.
+Our implementation takes approximately 15 minutes to reconstruct a single 2D slice on a Nvidia A100 GPU, so clinical adoption requires a more efficient implementation.
+Future research work will investigate the effectiveness of INRs as an image representation and explore the trade-offs between using diffusion models as regularization versus posterior sampling directly on the voxel grid.
 Acknowledgment
-This work was supported in part by NSF CCF-2239687 (CAREER), NSF IFML 2019844, and JCCO postdoctoral fellowship.
+Support in part by NSF CCF-2239687 (CAREER), NSF IFML 2019844, and JCCO postdoctoral fellowship.